OpenStudy (anonymous):

Pls help...evaluate 10^x when x = lg 5...thank you

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

I am assuming it is log base 10

OpenStudy (anonymous):

yes...is log base 10

OpenStudy (anonymous):

if you have base to the log of that base it cancel

OpenStudy (anonymous):

\[x^{\log_{x}b }=b\]

OpenStudy (anonymous):

what about if x = lg (lg 5)?

OpenStudy (anonymous):

still base 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You could just evaulate it using calculator

OpenStudy (anonymous):

gotta show workings

OpenStudy (anonymous):

5

OpenStudy (anonymous):

10^log[5]=ans log[10^log[5]]=log[ans] log[5]log[10]=log[ans] log[10]=1 log[5]=log[ans] get rid of log 5=answer

OpenStudy (anonymous):

i understood now but what about if x = lg (lg 5)?

OpenStudy (anonymous):

10^x=log[5]

OpenStudy (anonymous):

10^((10)^x)=5

OpenStudy (anonymous):

then use x=logy, where y=log5 10^x=y=log5

OpenStudy (anonymous):

did u get that kwoky

OpenStudy (anonymous):

im still trying to figure out ...

OpenStudy (anonymous):

ok, w/c one? the latter or the first q?

OpenStudy (anonymous):

the latter...cos i don't understand how can both answers turn out to be 5

OpenStudy (anonymous):

evaluate 10^x when x = log5 10^x=10^log5 = 5

OpenStudy (anonymous):

This one I understood but when x = log ( log 5), I don't quite understand.

OpenStudy (anonymous):

x = log ( log 5) then use x= log ( y ), where y=log5 10^x=10^logy =y =log5

OpenStudy (anonymous):

understood better?

OpenStudy (anonymous):

my pc kept hanging up/...lol

OpenStudy (anonymous):

erm....I think I'm quite dumb with log...still confused. I still think that with 10^lg ( lg 5), the 10^lg will be cancelled thus the ans should be lg 5 instead of 5.

OpenStudy (anonymous):

10^logy = y

OpenStudy (anonymous):

but since there is two log..thus shouldn't the 1st log be written off with the 10, therefore answer should left with 2nd log?

OpenStudy (anonymous):

therefore 10^log(log5) =log5

OpenStudy (anonymous):

yup...ans should be log 5 then.

OpenStudy (anonymous):

10^log(M) =M 10^log555 =555

OpenStudy (anonymous):

ok...thanks mark !

OpenStudy (anonymous):

ok wc kworky..good luck post some more q. so that you can practice solving them